OpenStudy (yacoub1993):

Solve the system of equations using matrices. Use Gaussian elimination with back-substitution. 3x + 5y - 2w = -13 2x + 7z - w = -1 4y + 3z + 3w = 1 -x + 2y + 4z = -5

OpenStudy (yacoub1993):

Anyone help me solve this one please

OpenStudy (da_scienceman):

Ok, first write down the system in augmented form AX=B where A are the coefficients of the system on d LHS, B is the vector on the RHS and X are the valiables. For example,

OpenStudy (yacoub1993):

3 5 -2 -13 2 7 -1 -1 4 3 3 1 -1 2 4 -5

OpenStudy (yacoub1993):

@da_ScienceMan I didnt understand how to do it. Could you please explain?

OpenStudy (da_scienceman):

Yes very well then....u almost got the matrix right. First let us correct the matrix

OpenStudy (yacoub1993):

ok

OpenStudy (yacoub1993):

then what is the second step @da_ScienceMan

OpenStudy (da_scienceman):

\[\left[\begin{matrix}3 & 5 & 0 & -2& | -13\\ 2& 0 & 7&-1 &|-1 \\0 & 4 & 3& 3 &| 1\\-1 & 2 & 4 & 0 & | -5\end{matrix}\right]\]

OpenStudy (da_scienceman):

this is the matrix A and then the vector is given by \[\left(\begin{matrix}x \\ y \\ z \\ w\end{matrix}\right)\]

OpenStudy (da_scienceman):

The next step is to reduce the matrix by guessing numbers such that u have almost all elements of the last rows equal to zero-

OpenStudy (da_scienceman):

Next, U proceed and have the second to the last row elements also zero but with 1 fewer than the last row....its a long process but u can do it.

OpenStudy (yacoub1993):

can you show me how to do it fro here. I am stuck over here exactly

OpenStudy (da_scienceman):

Yes sure hang on a min!

OpenStudy (da_scienceman):

i can tell u what to do step by step whil u write it down r u ready?

OpenStudy (da_scienceman):

first, I will divide row 2 by 2, next I multiply row 2 by 1 and add it to row 4. Then I have a new matrix.....

OpenStudy (yacoub1993):

Yes i am ready

OpenStudy (da_scienceman):

\left[\begin{matrix}3 & 5 & 0 & -2 & | -13 \\ 2 & 0 & 7 & -1 & | -1 \\0 & 4 & 3 & 3 & | +1 \\ -1 & 2 & 4 & 0 & | -5 \\\end{matrix}\right]rightarrow I will beginy by swapping rows 3 and row 4 to get the following matrix \[\left[\begin{matrix}3 & 5 & 0 & -2 & | -13 \\ 2 & 0 & 7 & -1 & | -1 \\-1 & 2 & 4 & 0 & | -5 \\ 0 & 4 & 3 & 3 & | 1 \\\end{matrix}\right]\] rightarrow

OpenStudy (da_scienceman):

i am trying to type the matrices after each step but it seems i made a mistake. Please swap row 3 and row 4

OpenStudy (da_scienceman):

Next divide row 2 by 2

OpenStudy (yacoub1993):

ok wait a minute

OpenStudy (da_scienceman):

\[\left[\begin{matrix}3 & 5 & 0 & -2 & | -13 \\ 2 & 0 & 7 & -1 & | -1 \\0 & 4 & 3 & 3 & | +1 \\ -1 & 2 & 4 & 0 & | -5 \\\end{matrix}\right]\rightarrow \left[\begin{matrix}3 & 5 & 0 & -2 & | -13 \\ 2 & 0 & 7 & -1 & | -1 \\-1 & 2 & 4 & 0 & | -5 \\ 0 & 4 & 3 & 3 & | 1 \\\end{matrix}\right]\]

OpenStudy (yacoub1993):

3 5 0 -2 | -13 1 0 3.5 -0.5 |-0.5 0 4 3 3 | 1 -1 2 4 0 |-5

OpenStudy (yacoub1993):

@da_ScienceMan

OpenStudy (da_scienceman):

then next step we get that \[\rightarrow \left[\begin{matrix}3 & 5 & 0 & -2 & | -13 \\ 1 & 0 & \frac{ 7 }{ 2 } & \frac{ -1 }{ 2} & | \frac{ -1 }{ 2 } \\-1 & 2 & 4 & 0 & | -5 \\ 0 & 4 & 3 & 3 & | 1 \\\end{matrix}\right]\]

OpenStudy (da_scienceman):

right tight. Next, U multiply row 2 by 1 and add it row 3!

OpenStudy (yacoub1993):

3 5 0 -2 |-13 1 0 7/2 -1/2 |-1/2 0 2 15/2 -1 | -11/2 0 4 3 3 | 1 @da_ScienceMan

OpenStudy (da_scienceman):

Very good. Next task: multiply row 3 by 2 and add it to row lets see what u get...remember our aim is to have the three first terms in row 4 all zero, first two terms in row 3 all zero and 1st term in row zero!

OpenStudy (da_scienceman):

multipliy row 3 by -2 and add it to row 4 i mean sorry!

OpenStudy (yacoub1993):

ok

OpenStudy (da_scienceman):

check the third term in row three in the above matrix u wrote and make SURE it is correct!

OpenStudy (yacoub1993):

3 5 0 -2 |-13 1 0 7/2 -1/2 |-1/2 0 -4 -15 2 | 11 0 0 -12 5 |12

OpenStudy (yacoub1993):

@da_ScienceMan is this correct

OpenStudy (da_scienceman):

ok I wrote the steps using my rough sheet and m sure u or me is wrong. Now let us go through it again if u dont mind...

OpenStudy (yacoub1993):

ok

OpenStudy (yacoub1993):

i made a mistake in the third row fourth term

OpenStudy (da_scienceman):

First u have to swap rows 3 and 4...which m sure u did. Next u multiply row 2 by 1/2 which u did.

OpenStudy (da_scienceman):

Sure u can just use my last matrix and proceed from there......I think it is correct! pls check properly its been a while I did this thing.

OpenStudy (yacoub1993):

ok wait let me write it down again

OpenStudy (da_scienceman):

Yes that would be nice...first swap rows3 and 4. Next multiply row 2 by 0.5. Then multiply row 2 by 1 and add to row 3. Next, multiply row 3 by -2 and add to row 4. Our aim is to have first three terms all zero in row 4, first two terms all zero in row and first term zero in row 3. Then we can proceed from there.

OpenStudy (yacoub1993):

@da_ScienceMan This is what i got 3 5 0 -2 |-13 1 0 7/2 -1/2 |-1/2 0 -4 -15 1 |-9 0 0 -12 4 |-8

OpenStudy (da_scienceman):

i didnt get that....d first two rows seem correct to me. but the last two???

OpenStudy (da_scienceman):

hang on a min. u just have to add d result u got from multiplying row 3 by adding them to row 4. keep row 3 as it is for nw!!!

OpenStudy (yacoub1993):

can you do it because i m confused

OpenStudy (da_scienceman):

so if u do it that way, u ll have [0 2 15/2 -1/2 | -11/2] for row 3, and [0 0 -12 4| 12] for row 4. pls verify!

OpenStudy (yacoub1993):

my last metrix is correct

OpenStudy (da_scienceman):

\[\left[\begin{matrix}3 & 5 & 0 & -2 & | -13 \\ 1 & 0 & \frac{ 7 }{ 2 } & \frac{ -1 }{2 } & \frac{ -1 }{ 2 } \\ 0 & 2 & \frac{ 15 }{ 2 } & \frac{ -1 }{2 } &| \frac{ -11 }{ 2 } \\ 0 & 0 & -12 & 4 &| 10\end{matrix}\right]\]

OpenStudy (da_scienceman):

is d expected matrix

OpenStudy (da_scienceman):

pls u can multiply row 3 and row 2 by 2 to make it easier for u to proceed!

OpenStudy (yacoub1993):

i am realy frustrated .. i will go out for dinner i will calculate later... can you just write down the steps and i will try solving it.. if i didnt understand we will continue later if you online

OpenStudy (yacoub1993):

@da_ScienceMan

OpenStudy (da_scienceman):

ok sure its just easy for sure. it is just because m kinda laid back typing d matrices. no worries if coming back i can write down d steps, nd makes them available too. Just relax nd go for dinner. I ll be on for sometime!

OpenStudy (yacoub1993):

OK THNXS ALOT @da_ScienceMan

OpenStudy (da_scienceman):

sure no worries!

OpenStudy (da_scienceman):

please check the steps in d attached file and verify. I represented d steps using R for rows together with the corresponding operation....

OpenStudy (da_scienceman):

I claim that x is 6.6, y is -1.4, z is 6.3 and w is 1.1. But please verify!